OpenStudy (anonymous):

The length of the top of a table is 7m greater than the width. The area is 78m^2. Find the dimensions of the table

OpenStudy (zzr0ck3r):

length*width = area lets say l*w = a l = w+7 per the first line of your question l*w = 78 by the second line of your question plug w+y for l and solve